OpenStudy (anonymous):

In writing the ratio of measurement numbers, the units must be the same. True False

OpenStudy (anonymous):

Think with me... A ratio is a unitless number... So now give me the answer

OpenStudy (anonymous):

False?

OpenStudy (anonymous):

how do you get a unitless number? let's say 30 Newtons and 15 Newtons if you devide both you'll have \[\frac{ 30 }{ 15 } \times \frac{ Newton }{ Newton } = 2\] unitless

OpenStudy (anonymous):

I kinda still don't get it.

OpenStudy (anonymous):

When you devide 2 things that are the same the answer is 1 so when you devide Newton by Newton (force measure) you'll get 1 That's why 30 N / 15 N = (30/15) * (N/N) , where (N/N) = 1 so the answer is just 30/15 = 2